OpenStudy (anonymous):

1.4x+3.9y=6.4 0.2x-1.3y=1.2 solve using substitution method

OpenStudy (anonymous):

HI :DDD

OpenStudy (anonymous):

The substitution method asks you to find what either x or y equals, and to then take that whole equality and substitute if for either your x or your y in the other equation. In this case, since I loathe decimals, I'm going to scale each equation by a factor of ten to give: 14x + 39y = 64 2x - 13y = 12 The easiest to find a substitution for is x, using the second equation. We can see that 2x - 13y = 12 can be rewritten as 2x = 13y + 12, and so x = 13/2 y +6. Thus we substitute x = 13/2 y + 6 into our first equation, to give: 14(13/2 y + 6) + 39y = 64 I'll leave it there since from thereon this is just simplifying and solving, but let me know if you'd like the completed solution.